Find the (a) mean, (b) median, (c) mode, and (d) midrange for the data and then (e) answer the given question Listed below are the weights in pounds of 11 players randomly selected from the roster of a championship sports team. Are the results likely to be representative of all players in that sports league? 260 270 180 187 303 210 298 309 180 8 298 a. Find the mean. The mean is pound(s). (Type an integer or a decimal rounded to one decimal place as needed.) b. Find the median. The median is pound(s). Type an integer or a decimal rounded to one decimal place as needed.) c. Find the mode. Select the correct choice below and, if necessary, fill in the answer box to complete your choice. O A. The mode(s) is(are) pound(s) O B. There is no mode. d. Find the midrange. The midrange is pound(s). (Type an integer or a decimal rounded to one decimal place as needed.)
Answer #1

a) Mean = (260 + 270 + 180 + 187 + 303 + 210 + 298 + 309 + 180 + 181 + 298)/11

= 243.3 pounds

b) Data in ascending order : 180, 180, 181, 187, 210, 260, 270, 298, 298, 303, 309

There are 11 terms

Median is (n+1)/2 th term = (11+1)/2 = 6th term

Median = 260 pounds

c) A. The modes are 180, 298 pounds

d) Mid range = (180 + 309)/2

= 244.5 pounds

e) The measures of central tendencies are not approximately equal. The sample is not representative of population. The sample is drawn from a cluster within the population. So, we cannot say that the sample is representative of all the players in that sports league